The pH of water at 50 degrees C is 6.55. At this temperature, the hydronium ion concentration is ___________ mol/L and the value of Kw is ______________:
Answer #1

H2O = H3O+ + OH-

pH = - log [H3O +] = 6.55

So, [H3O +] =10^-6.55 =2.82*10^-7 mol/L

Kw = [H3O +][OH-] = (2.82*10^-7)^2 =7.94*!0^-14